Evaluate -2-(-6+(-3))
Answer by lenny460(775)

-2 -(-6 + (-3))
-2 -1(-6 + -3)
Combine like terms:
-6 + -3 = -9
-2 -1(-9)
Multiply:
-1 * -9 = 9
-2 + 9 = 7
The Answer:
7


obuong3@aol.com